OpenStudy (anonymous):

Use complete sentences to analyze the graph of the function f(x) = log (base 1/4) x descrive teh domain, range and general shape of the graph

OpenStudy (anonymous):

try to find the inverse function of it. then look of its Domain and Range. Then (D->R,R->D) (f^-1->f) Then the origenal function is a reflection along y=x. Then you can see how it looks.

OpenStudy (anonymous):

i dont think thats right....

OpenStudy (anonymous):

Do you how the inverse function is relating to a function?

OpenStudy (phi):

I would plot a few key points to get the general shape. \( log_{\frac{1}{4}}(x) =y \) corresponds to \( (\frac{1}{4})^y = x \) we know that no matter what power we choose 1/4 will not be negative, so x>0 \( (\frac{1}{4})^{1} = \frac{1}{4} \) so \( log_{\frac{1}{4}}(\frac{1}{4}) =1 \) \( (\frac{1}{4})^{0} = 1 \) so \( log_{\frac{1}{4}}(1) =0 \) \( (\frac{1}{4})^{-1} = 4 \) so \( log_{\frac{1}{4}}(4) = -1 \) sketch the graph you can see the domain is (0, inf) and though it is slowly going down, it can reach any negative number, so its range is (-inf, inf)